OpenStudy (anonymous):

Car headlights have both low beams and high beams. The high beams may be necessary when it is very dark outside. If the power for the high beam is 60.0 watts and the current is 5.0 amps, how much voltage is required?

OpenStudy (dape):

The power P in a circuit with current I and voltage U is simply related by \[ P = U*I \] Plugging in the givens we get \[ (60 W) = U*(5 A) \] So U = 12 V